# Skyrme-Hartree-Fock-Bogoliubov mass models on a 3D mesh: IV. Improved description of the isospin dependence of pairing

### v1

@@ 3.2 Fission properties of BSkG4
In addition to the rms deviations shown in Table 2, this section analyzes the differences between BSkG3 and BSkG4 in terms of fission properties. Fig. 7 displays the primary and secondary fission barriers (upper and middle panels, respectively) as well as the isomer excitation energy (lower panel) for the $Z\geq 90$ nuclei with known empirical barriers included in the RIPL-3 database Capote09 . The numerical details of the calculations performed with the new model are the same as those in the previous work on BSkG2 Ryssens23 .
In the upper panel of Fig. 7 we see that BSkG4 predicts higher primary barriers for most nuclei, compared to BSkG3. As a result, the new parameterization is in closer alignment with the empirical values, especially for the heaviest isotopes of the chains shown (e.g., 240-245Pu, 243-244Am, and 241-249Cm). This improvement is directly attributable to the new pairing treatment: BSkG3 predicted lower primary barriers due to an overestimation of pairing effects.
A similar pattern is found for the secondary barriers (middle panel of Fig.7). With the exception of the Th and Pa chains, BSkG4 generally produces higher values than BSkG3, with these differences becoming more pronounced as the neutron number increases. Although in some isotopic chains (e.g., Pu) BSkG4 predictions deviate more from the reference values compared to BSkG3, the overall level of agreement remains comparable (see rms deviations in Table 2).

### v2

@@ 3.2 Fission properties of BSkG4
In addition to the mean and rms deviations shown in Table 2, this section analyzes the differences between BSkG3 and BSkG4 in terms of fission properties. Fig. 7 displays the primary and secondary fission barriers (upper and middle panels, respectively) as well as the isomer excitation energy (lower panel) for the $Z\geq 90$ nuclei with known empirical barriers included in the RIPL-3 database Capote09 . The numerical details of the calculations performed with the new model are the same as those in the previous work on BSkG2 Ryssens23 .
In the upper panel of Fig. 7 we see that BSkG4 predicts slightly higher primary barriers for most nuclei, compared to BSkG3. This is directly attributable to the new pairing treatment: BSkG3 predicted lower primary barriers due to an overestimation of pairing effects. The differences become larger as the mass increases. for instance, Am and Cm chains present more deviations in comparison with Pa or U.
A similar pattern is found for the secondary barriers (middle panel of Fig.7). With the exception of the Th chain, BSkG4 generally produces higher values than BSkG3, with these differences becoming more pronounced as the neutron number increases. Although in some isotopic chains (e.g., Th or Cm) BSkG4 predictions deviate more from the reference values compared to BSkG3, the overall level of agreement remains comparable (see rms deviations in Table 2).